tml Help Save & Exlt Submit 2 e of rmu materials 290,000 s 100.000 s 285,000o 270,000 nInventory balances at the beginning and end of the year were as follows of rear tad。E Year s 40,000 10,000 2 35,000 Hork in process goods s 50,000 The total manufacturing costs for the year were $683,000; the cost of goods available for sale of goods sold totaled $660,000; and the net operating income was $30,000. The companys underapplied or totaled $740.000; the unadjusted cost Prepare schedules of cost of goods manufactured and cost of goods sold and an income statement. (Hint Prepare of goods sold first the schedule of cost of goods manufactured.) Prey 2 of 3EENext: > MacBook Air
Answer #1
Superior Company
Income Statement
Sales revenue 915000
Cost of goods sold 645000
Gross profit 270000
Selling and administrative expenses:
Selling expenses 140000
Administrative expenses 100000
240000
Net operating income 30000
Superior Company
Schedule of Cost of Goods Sold
Cost of goods manufactured 690000
Add: Beginning finished goods inventory 50000
Cost of goods available for sale 740000
Less: Ending finished goods inventory 80000
Unadjusted cost of goods sold 660000
Less: Overapplied manufacturing overhead (285000 - 270000) 15000
Adjusted cost of goods sold 645000
Superior Company
Schedule of Cost of Goods Manufactured
Direct materials:
Beginning raw materials inventory 40000
Add: Purchases of raw materials 290000
Total raw materials available 330000
Less: Ending raw materials inventory 10000
Raw materials used in production 320000
Direct labor 78000
Manufacturing overheads 285000
Total manufacturing costs for the year 683000
Add: Beginning work in process inventory 42000
Total work in process 725000
Less: Ending work in process inventory 35000
Cost of goods manufactured 690000

Working:

A B C D
Superior Company
Income Statement
4 Sales revenue =D5+D6
5 Cost of goods sold 645000
6 Gross profit =D11+D12
7 Selling and administrative expenses:
8 Selling expenses 140000
9 Administrative expenses 100000
10
11 =SUM(C8:C10)
12 Net operating income 30000
Superior Company
Schedule of Cost of Goods Sold
16 Cost of goods manufactured =C18-C17
17 Add: Beginning finished goods inventory 50000
18 Cost of goods available for sale 740000
19 Less: Ending finished goods inventory =C18-C20
20 Unadjusted cost of goods sold 660000
21 Less: Overapplied manufacturing overhead (285000 - 270000) =285000-270000
22 Adjusted cost of goods sold =C20-C21
Superior Company
Schedule of Cost of Goods Manufactured
Direct materials:
27 Beginning raw materials inventory 40000
28 Add: Purchases of raw materials 290000
29 Total raw materials available =C27+C28
30 Less: Ending raw materials inventory 10000
31 Raw materials used in production =C29-C30
32 Direct labor =D34-D31-D33
33 Manufacturing overheads 285000
34 Total manufacturing costs for the year 683000
35 Add: Beginning work in process inventory =D36-D34
36 Total work in process =D38+D37
37 Less: Ending work in process inventory 35000
38 Cost of goods manufactured 690000